OpenStudy (anonymous):

Medal a = -2 , b=4, c=-1 , d=3 -3/4 b + 1/2 (ac+bd)

OpenStudy (dan815):

when you seet the letters change that with the numbers they gave you

OpenStudy (anonymous):

-3/4 *4 + 1/2 (-2*-1+4*3)=4

OpenStudy (anonymous):

-3/4 (4) + 1/2 (-2) (-1) + (4) (3)

OpenStudy (dan815):

-3/4 * 4 +1/2 (-2*-1+bd) can you replace the bd part yourself now

OpenStudy (anonymous):

i already did it haha

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the answer is 4

OpenStudy (anonymous):

how u get this

OpenStudy (anonymous):

because all you have to do iss replace the letters with numbers they give you

OpenStudy (anonymous):

do you need help with more?

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

can you show me how did u get it

OpenStudy (anonymous):

-3/4 b + 1/2 (ac+bd)

OpenStudy (anonymous):

they tell you b is 4

OpenStudy (anonymous):

so remove b and put 4 instead

OpenStudy (anonymous):

and they tell you a is -2 so replace a with -2

OpenStudy (anonymous):

how would you solve it

OpenStudy (anonymous):

whenyou replace all letterers with numbers 3/4(4)+1/2(−2*−1+4*3) =−3+1/2(−2*−1+4*3) =−3+1/2(2+4*3) =−3+1/2(2+12) =−3+1/2(14) =−3+7 =4

OpenStudy (anonymous):

get it now?

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

you're welcome need help with more?

OpenStudy (anonymous):

thanks for the medal :)